Boxcars, Inc. has a capital structure consisting of $100MM in equity and $150MM in debt. With this capital structure, the expected return to its equity is 18 percent, the expected return to its debt is 7 percent, and the market beta of Boxcars enterprise value is 1.68. The risk-free rate is 3 percent. The firm unexpectedly issues $110MM in new shares, using the cash to repurchase $110MM of its debt. After the repurchase, the remaining $40MM in debt is risk-free because there is no chance the firm will default on the debt.

a. What is the Boxcar’s weighted average cost of capital prior to the change in capital structure?
Answer #1

debt/equity = 150/100 =1.5

D/A = D/(E+D)
D/A = 1.5/(1+1.5)
=0.6
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 7*(1-0)
= 7
Weight of equity = 1-D/A
Weight of equity = 1-0.6
W(E)=0.4
Weight of debt = D/A
Weight of debt = 0.6
W(D)=0.6
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=7*0.6+18*0.4
WACC% = 11.4